OpenStudy (anonymous):

find me that limit please: limit x*ln^3(x) as x->0+ show the used method if possible!

OpenStudy (experimentx):

0??

OpenStudy (experimentx):

is that (ln x)^3 ??

OpenStudy (experimentx):

use L'hospital rule!! put x = 1/(1/x)

OpenStudy (anonymous):

but still wont get rid of the power!!

OpenStudy (experimentx):

try 3 times

OpenStudy (anonymous):

what about using nth root since we have 3 as a power!

OpenStudy (experimentx):

since it is a limit ... i think it should work out fine ... better elegant than the method i suggested \[ \left ( \lim_{x->0}\frac{\ln x}{(1/x)^{1/n}}\right )^n\]

OpenStudy (experimentx):

Oh ... sorry ... that n is 3

OpenStudy (anonymous):

yep i know thank's :D